Status Report

NASA Space Station On-Orbit Status 20 January 2005

By SpaceRef Editor
January 20, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 20 January 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Preparations continued for the Orlan EVA-12 on 1/26:

At the battery charger (ZU-S) in the DC1 docking module, FE Salizhan Sharipov terminated charging on the second 28V battery pack (825M3) for the Orlan backpack and removed the pack from the charger.

Other EVA preparations today included: (1) checkout of the Orlan BSS water/gas separation systems at the Service Module Transfer Compartment (SM PkhO) and in the DC1 docking module, (2) gathering and servicing the Orlan suit consumable/replaceable ORU elements and personal gear, (3) activating and inspecting the two Orlan spacesuits #25 and #27, (4) performing KVD pressure equalization valve functionality tests on the suits and their BSS interface units in the DC1 and SM PkhO (backup). Orlan prepping activities were supported by peregovoriy (tagup) with ground spets (specialist). [Orlan ORUs are LiOH canisters (LP-9), primary & backup oxygen tanks (BK-3), moisture collectors, feedwater filters (FOR), CO2 measuring unit (IK) filter, filtration & separation units (BOS), and the newly charged 825M3 storage batteries.]

The crew installed a US EMU (extravehicular mobility unit) helmet light on Orlan #27 and checked out the already installed (from Increment 9) EMU light on Orlan #25, and also removed the OTA (Orlan tether adapter) from the old #14 suit and mounted it on #27.

The video imagery of the EVA equipment and tools taken yesterday was downlinked and discussed with ground specialists via tagup. [The footage included views of the URM-D baseplate with platform, the TM/TS “monoblock”, the ROBOTIK monoblock, the packed EVA tool carrier (KPU), the BIORISK (MSN) package, a general view of configured EVA bundles #1 and #2, and locations for pry bar and power screwdriver.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Sharipov completed the periodic (weekly) replenishing of the Elektron’s water supply for electrolysis, his eleventh, filling the KOV thermal loops EDV container (#1013) with purified (deionized) water from the BKO multifiltration/purification column. [The procedure is specially designed to prevent air bubbles from getting into the BZh liquid unit where their pressure spikes, from collapsing, could cause micropump impeller cavitation and Elektron shutdown, as happened numerous times in the past. In the procedure, the EDV water is drawn from the BKO and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number).]

CDR/SO Leroy Chiao conducted the weekly data take with the two newly delivered CSA-CP (Compound Specific Analyzer-Combustion Products) units #1018 and #1019, which monitors the ongoing decontamination (outgassing) of the deployed instruments (last time done: 1/12). [CSA-CP measures O2 (oxygen), CO (carbon monoxide), HCN (hydrogen cyanide), and HCl (hydrogen chloride).]

Salizhan did the routine maintenance of the SM’s SOZh environment control & life support system, today including the weekly inspection of the BRPK air/liquid condensate separator apparatus.

The FE also installed an IP-1 airflow sensor in the SM transfer tunnel to the Progress 16 and completed the routine checkup of the IP-1s in the various RS (Russian segment) hatchways and FGB-to-Node tunnel.

Working off the Russian task list, Salizhan prepared the regular IMS (inventory management system) delta file for export/import to the IMS databases.

The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser and VELO cycle with bungee cord load trainer. Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 2 of a new set) and one hour on VELO.

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

Predicted times of station entry and exit into radiation zones from the recent energetic SPE (solar proton event) were uplinked to the crew for today.

Yesterday’s C&C (Command & Control) computer transition went well. [The system is now operating with C&C-2 as Primary, C&C-1 as Backup, and C&C-3 in Standby. The crew was thanked again for replacing the PCS (portable computer system) hard drive in the SM PCS, which became necessary after the reconnection of the PCS machines following the swap, and also for their continued support of these ground activities.]

The deployment of the new IBM A31p ThinkPad NGLs (next generation laptops) in the RS has now been certified by Safety experts on both sides, and the activity has been added to the “job jar” task list.

This week, the GN&C (Guidance, Navigation & Control) system is undergoing two major activities to enhance and better understand system performance: (1) tuning of SIGI (Space Integrated Global Positioning System [GPS] Inertial Navigation System [INS]), and (2) the CMG-2 (Control Moment Gyro #2) wheel speed test. [Late last year, controllers loaded a new firmware version (R2) to SIGI-2. After the load, the selection filter rejected the attitudes produced by SIGI-2. Through a series of PPLs (pre-positioned loads), the ground is attempting to fine-tune the SIGI-2 to get better reception and to track good attitudes. The rotor speed test will increase the baseline vibration and spin motor current information for the CMGs. Both test activities will conclude on Friday.]

At ~11:40am EST, the ground uplinked a software patch to the Lab PCS laptop that will eliminate the need for the BGA (Beta gimbal assembly) navigation workaround. The crew was to reboot the PCS if necessary. If all went well with the Lab PCS, the rest of the PCSs were to be loaded.

Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Tsunami Damage, Mauritius (Dynamic Event. The island of Mauritius reports some damage from the recent tsunami, and portions of the eastern coastline were completely submerged by the wave. Looking to the left of track for regions of standing water inland and bare soil along the eastern shoreline), Tsunami Damage, Maldives (Dynamic Event. This nadir pass provided an opportunity for photography of the southern Maldives. Looking for regions for standing water inland, and bare soil along the coastlines. Standing bodies of salt water inland may act as contaminant point sources to the islands’ fresh water aquifers), Sobat fans, SE Sudan (this nadir overpass provided an opportunity for mapping swaths across the central portion of this megafan complex. A highly complex series of individual fans are present, and high-resolution imagery will help delineate the fan boundaries. Looking for divergent drainage patterns and wetland boundaries to locate fans), Patagonian Glaciers, S. America (this nadir pass provided an opportunity for mapping ice extent in the northern glacier fields. Mapping swaths across the leading edges of the smaller mountain glaciers are useful for assessment of ice volume and motion), and Pilcomayo River dynamics, N. Argentina (weather is predicted to be mostly clear over the river. Mapping of prior and current river courses is useful for understanding the dynamics of the river system. Sunglint may help highlight smaller tributary channels).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • EVA-12 — 1/26/05 (hatch opening 2:27am EST)
  • Progress 16P undocking & destructive reentry — 2/27/05;
  • Progress 17P launch — 2/28/05.
  • EVA-13 — 3/25/05;
  • Soyuz 10 S launch — 4/15/05;
  • Soyuz 9S undock — 4/25/05 (after 193 days on orbit, 191 days on board ISS).

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 6:22am EST [= epoch]):

  • Mean altitude — 358.7 km
  • Apogee height — 365.0 km
  • Perigee height — 352.3 km
  • Period — 91.72 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.000942
  • Solar Beta Angle — -17.3 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 70
  • Revolutions since FGB/Zarya launch (Nov. 98) — 35245

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.